OpenStudy (sidonce):

how many molecules has 500 ml 0.05 M of H2SO4 solution?

OpenStudy (pnathaniel56):

i dont know

OpenStudy (pnathaniel56):

its taking u so long

OpenStudy (pnathaniel56):

its taking u so long

OpenStudy (sapphiremoon):

Okay, molarity is measured in terms of mol/L, so all we need is some stoichiometry to figure this out. We have milliliters, and we want moles. We'll convert mL to L, so we have 0.5 L. With the liters and moles per liter we can find the moles like so:\[0.5 L \times \frac{ 0.05 mol }{ L } = 0.025 mol\] because the liters cancel. Does this make sense?

OpenStudy (sidonce):

n=C*V=0,05*0,5=0,025mol N=n*Na=0,025*6,02*10^23=0,15*10^23 <-- the answer, i quess so but i dont know is it's right and if the formules are right chosed

OpenStudy (sapphiremoon):

Yes, you're correct. Multiply the moles by Avogadro's number, and you should get a rather large number of molecules. Remember that in scientific notation, it's not 0.15x10^23, it's 1.5x10^22.

OpenStudy (sidonce):

thank you :)

OpenStudy (sapphiremoon):

You're very welcome! :)

OpenStudy (cuanchi):

how many molecules has 500 ml 0.05 M of H2SO4 solution? the question is wrong formulated. It asked about molecules, but it doesnt said molecules of what? If it is a solution in water you can be able to consider also the molecules of water and they are much more than the molecules of sulfuric acid. By the way the sulfuric acid will be hydrolyzed and will be not molecules of sulfuric acid in the solution but polyatomic anions and cations